OpenStudy (anonymous):

complete the ordered pair for the equation y=-9x+3 ( ,30)

OpenStudy (anonymous):

that is the y value so lets substitute it for y -7 = 2x + 5 -12 = 2x -6 = x

OpenStudy (anonymous):

so -6

OpenStudy (anonymous):

well to complete it it would be (-6, -7)

OpenStudy (anonymous):

what aboit the one i just posted